OpenStudy (anonymous):

subtract 8x/x+1 - 3/x+1 please help ?

OpenStudy (anonymous):

Is this like:\[\frac{ 8x }{ x+1 } - \frac{ 3 }{ x+1 }\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok, well, it looks like lots of "stuff", but if you think about when you were just doing basic subtraction or addition with normal (non-algebra) fractions, if they had a common denominator, you could add or subtract directly... like 4/7 + 1/7 = 5/7

OpenStudy (anonymous):

here your common denominator is x+1 But you can still add the stuff on top, and combine it all over the common denominator, x+1

OpenStudy (anonymous):

(well, it's subtracting here, not adding, but same idea)

OpenStudy (anonymous):

\[\frac{ 8x }{ x+1 } - \frac{ 3 }{ x+1 } = \frac{ (8x-3) }{ x+1 }\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you are welcome... glad to help! Hope that made sense!